# UTG: Towards a Unified View of Snapshot and Event Based Models for Temporal Graphs

### v1

@@ 6.4 UTG Training Ablation Study
Table 4: GCLSTM ablation with UTG training when compared with Pytorch Geometric-Temporal training, results averaged across five runs best results are bolded.
Dataset GCLSTM (UTG) GCLSTM (original)
UCI 0.093 $\pm$ 0.006 0.047 $\pm$ 0.006
Enron 0.170 $\pm$ 0.006 0.131 $\pm$ 0.003
Contacts 0.128 $\pm$ 0.004 0.101 $\pm$ 0.031
Social Evo. 0.286 $\pm$ 0.003 0.287 $\pm$ 0.009
MOOC 0.143 $\pm$ 0.006 0.076 $\pm$ 0.003
Table 4 compares the performance of GCLSTM using UTG training with that of Pytorch Geometric Temporal training on the DTDG datasets (Section 5.2). UTG training significantly enhances GCLSTM performance across four datasets and performs identically on the Social Evo dataset. This demonstrates the benefits of predicting the next snapshot and backpropagating after every snapshot.
@@ 7 Conclusion
Appendix A Broader Impact
We expect our proposed UTG framework to impact the field of temporal graph learning in a positive way by allowing unified comparison between previously disjoint lines of research. The empirical findings from this paper suggest the potential to combine the advantages from both snapshot and event-based methods to design effective and scalable temporal graph methods in the future. The UTG framework can also be applied to any existing and future snapshot or event-based methods to enable them to evaluate on both types of temporal graph representations.
Appendix B Potential Negative Impact
Our proposed UTG is a framework which unifies different types of methods from temporal graph learning thus there are no direct negative impacts. However, temporal graph learning can be widely applied in applications such as fraud detection, traffic forecasting, recommendation systems and more. Without proper understanding of the behavior of TG methods, it is possible that the model predictions might be incorrect or misleading thus causing negative impact.
Appendix C Limitations
In this work, we focus on the streaming setting, which is the most commonly used setting for evaluation of event-based methods. However, other settings such as the deployed setting and the live-update setting as discussed in Appendix E are not considered due to computational constraints. Future work can utilize the UTG framework under the aforementioned settings.
Appendix D Computational Time
@@
Table D shows the inference time for all methods on DTDG datasets. Table D shows the inference time for all methods on CTDG datasets. OOM means out of memory and OOT means out of time. We observe that snapshot-based models are at least one order of magnitude faster than event-based models such as NAT, DyGFormer and GraphMixer. In addition, the best performing model on most datasets, DyGFormer, is also consistently the slowest method.
Appendix E Evaluation Settings
@@
Deployed setting. The deployed setting is often used as the evaluation setting for snapshot-based methods [30, 15]. In this setting, no information from the test set is passed to the model, and the node embeddings from the last training snapshot are used for predictions in all test snapshots.
Streaming setting. event-based models often evaluate with the streaming setting [18, 33, 26, 12]. In this setting, information from the previously observed batches of events can be used to update the model however no information from the test set can be used to train the model.
Live-update Setting. You et al. [32] proposed the live-update setting where the model weights are constantly updated to newly observed snapshots while predicting the next snapshot. To predict links in $\mathbf{G}_{t+1}$, first the observed snapshot $\mathbf{G}_{t-1}$ are split into a training set and a validation set. The model is trained on $\mathbf{G}_{t-1}^{train}$ while using $\mathbf{G}_{t-1}^{val}$ for early stopping. Lastly, the trained model receives $\mathbf{G}_{t}$ and predicts for $\mathbf{G}_{t+1}$.
@@
Appendix F Temporal Graph Learning Methods
F.1 Snapshot-based Methods
ROLAND [32]: You et al. [32] propose ROLAND as a framework for repurposing static GNNs to DTDGs. In ROLAND, a hierarchy of node states is constructed by different GNN layers that are recurrently updated over time. You et al. [32] introduce a novel evaluation setting, live-update, where GNNs are updated based on the recently observed events after making prediction. ROLAND provides the opportunity of adapting efficient static GNN designs for DTDGs.
HTGN [30]: HTGN utilizes hyperbolic geometry to better capture the complex and hierarchical nature of the evolving networks. It combines hyperbolic GNNs and hyperbolic gated recurrent network to jointly model temporal dynamics and hierarchical structures. HTGN consists of two main modules: hyperbolic temporal contextual self-attention that focuses on the historical state attention, and hyperbolic temporal consistency that provides stability and generalization.
GCLSTM [2]: GCLSTM learns the node structure s through a GCN, while the temporal features are captured with an LSTM. The models is proposed especially for dynamic link prediction task.
EGCN [15]: EvolveGCN focuses on dynamic graph representation learning for temporal networks where the set of nodes changes frequently. Basically, EvolveGCN exploits an RNN to adapt the parameters of a GCN over time. By focusing on adapting the model instead of the node embeddings, EvolveGCN is not restricted to the node set and can learn representation for the nodes without historical information.
PyG-Temporal [20]: PyTorch Geometric Temporal is a comprehensive framework that facilitate neural spatiotemporal signal processing. It offers a cohesive platform that simplifies temporal geometric deep learning. PyG Temporal is built upon the foundations of existing PyTorch libraries, and it includes streamlined neural network layer definitions, efficient temporal snapshot generators for batching, and a variety of benchmark datasets.
F.2 Event-based Methods
TGAT [28]: It is one of the pioneering work on inductive representation learning on temporal graphs. Xu et al. [28] argue that learning on dynamic graphs requires methods that handle new nodes and capture temporal patterns, with node embeddings representing both static features and evolving structures. The temporal graph attention (TGAT) layer is proposed to efficiently aggregate temporal-topological neighborhood features and learn time-feature interactions using self-attention. Stacking TGAT layers enables the model to treat node embeddings as functions of time, allowing inductive inference for new and existing nodes.
@@ F.2 Event-based Methods
TGN [18]: Rossi et al. [18] introduce Temporal Graph Networks (TGNs), a versatile and efficient framework for dynamic graphs, represented as stream of timestamped events. TGNs leverage a combination of memory modules and graph-based operators to improve computational efficiency. Essentially, TGN is a framework that subsumes several previous models as specific instances.
@@
CAWN [26]: To learn the underlying laws, such as triadic closure, that governs real-world network, Wang et al. [26] propose Causal Anonymous Walks (CAWs) for representing temporal networks inductively. CAWs, derived from temporal random walks, act as automatic retrievals of temporal network motifs, avoiding the need for their manual selection and counting. Wang et al. [26] alswo propose an anonymization strategy which replaces node identities with hitting counts from sampled walks, maintaining inductiveness and motif correlation. CAWN is a neural network model proposed to encode CAWs, paired with a CAW sampling strategy that ensures constant memory and time costs for online training and inference.
@@
TCL [25]: TCL effectively learns dynamic node representations by capturing both temporal and topological information. It features three main components: a graph-topology-aware transformer adapted from the vanilla Transformer, a two-stream encoder that independently extracts temporal neighborhood representations of interacting nodes and models their interdependencies using a co-attentional transformer, and an optimization strategy inspired by contrastive learning. This strategy maximizes mutual information between predictive representations of future interaction nodes, enhancing robustness to noise.
@@
NAT [12]: In modeling temporal networks, the neighborhood of nodes provides essential structural information for interaction prediction. It is often challenging to extract this information efficiently. Luo and Li [12] propose the Neighborhood-Aware Temporal (NAT) network model that introduces a dictionary-type neighborhood representation for each node. NAT records a down-sampled set of neighboring nodes as keys, enabling fast construction of structural features for joint neighborhoods. A specialized data structure called N-cache is designed to facilitate parallel access and updates on GPUs.
EdgeBank [16]: Poursafaei et al. [16] introduce stricter evaluation procedures for dynamic graph link prediction, reflecting real-world scenarios. They leverage the recurring patterns of edges to proposes EdgeBank that is a baseline relying on memorization with surprisingly good performance.
TGB [7]: This is a curated collection of comprehensive and varied benchmark datasets tailored for rigorous evaluation of machine learning models on temporal graphs. These datasets are extensive and diverse and the benchmark includes tasks for both node-level and edge-level prediction, with evaluation protocols designed around practical use cases. TGB provides an automated machine learning pipeline for reproducible and accessible research, encompassing data loading, experiment setup, and performance evaluation.
@@
DyGFormer [33]: Yu et al. [33] introduces a transformer-based architecture for dynamic graph learning. DyGFormer focuses on learning from nodes historical first-hop interactions and employs a neighbor co-occurrence encoding scheme to capture correlations between source and destination nodes through their historical sequences. Yu et al. [33] also propose a patching technique that divides each sequence into patches for the transformer, enabling effective utilization of longer histories. Yu et al. [33] also present DyGLib that is a library for standardizing training pipelines, extensible coding interfaces, and thorough evaluation protocols to ensure reproducible dynamic graph learning research.
Appendix G Computing Resources
@@
Appendix H Model Configurations

### v2

@@ 5 Experiments
Table 4: Base models with UTG training when compared to original training, results averaged across five runs, best results for each model are bolded.
Dataset GCLSTM (UTG) GCLSTM (original) EGCNo (UTG) EGCNo (original) HTGN (UTG) HTGN (original)
UCI 0.093 $\pm$ 0.006 0.047 $\pm$ 0.006 0.121 $\pm$ 0.010 0.124 $\pm$ 0.007 0.093 $\pm$ 0.012 0.052 $\pm$ 0.006
Enron 0.170 $\pm$ 0.006 0.131 $\pm$ 0.003 0.233 $\pm$ 0.008 0.227 $\pm$ 0.012 0.267 $\pm$ 0.007 0.196 $\pm$ 0.025
Contacts 0.128 $\pm$ 0.004 0.101 $\pm$ 0.031 0.192 $\pm$ 0.001 0.195 $\pm$ 0.001 0.165 $\pm$ 0.001 0.129 $\pm$ 0.020
Social Evo. 0.286 $\pm$ 0.003 0.287 $\pm$ 0.009 0.253 $\pm$ 0.006 0.231 $\pm$ 0.009 0.228 $\pm$ 0.003 0.178 $\pm$ 0.024
MOOC 0.143 $\pm$ 0.006 0.076 $\pm$ 0.003 0.126 $\pm$ 0.009 0.119 $\pm$ 0.009 0.093 $\pm$ 0.005 0.079 $\pm$ 0.011
Benefits of UTG Training. UTG training enables snapshot-based models to effectively incorporate novel information during inference. Table 4 shows the performance benefit of using UTG training on the GCLSTM, HTGN and EGCNo models when compared to the original training scheme (such as in Pytorch Geometric Temporal) for DTDG datasets. UTG training significantly enhances model performance across most datasets and performs identically on the rest.
Discussion. Event-based methods such as NAT and DyGFormer tend to perform best on both CTDG and DTDG, potentially leading to the premature conclusion that event-based modeling is the preferred paradigm and that snapshot-based models should be avoided. However, the superior performance of NAT and DyGFormer could be primarily due to their ability to leverage joint neighborhood structural features, specifically the common neighbors between the source and destination nodes of a link [17, 28]. This approach has been shown to be fundamental for achieving competitive link prediction on static graphs [38, 39]. In contrast, none of the existing snapshot-based methods incorporate joint neighborhood structural features. Therefore, the performance difference could be mainly attributed to this factor rather than an intrinsic difference between event-based and snapshot-based models. This is suggested by the fact that event-based models such as TGN and GraphMixer, which omits these features, have no clear performance advantage over snapshot-based methods. Moreover, snapshot-based methods are more computationally efficient and might be preferred when efficiency is important. These considerations suggest that both event-based and snapshot-based methods have their own merit. We believe that combining the strengths of both approaches is an important future direction.
@@ 6 Conclusion
Acknowledgements
This research was supported by the Canadian Institute for Advanced Research (CIFAR AI chair program), Natural Sciences and Engineering Research Council of Canada (NSERC) Postgraduate Scholarship Doctoral (PGS D) Award and Fonds de recherche du Québec - Nature et Technologies (FRQNT) Doctoral Award.
Appendix A Computational Time
@@
Table A shows the inference time for all methods on DTDG datasets. Table A shows the inference time for all methods on CTDG datasets. OOM means out of memory and OOT means out of time. We observe that snapshot-based models are at least one order of magnitude faster than event-based models such as NAT, DyGFormer and GraphMixer. In addition, the best performing model on most datasets, DyGFormer, is also consistently the slowest method.
Appendix B Evaluation Settings
@@
Deployed setting The deployed setting is often used as the evaluation setting for snapshot-based methods [12, 14]. In this setting, no information from the test set is passed to the model, and the node embeddings from the last training snapshot are used for predictions in all test snapshots.
Streaming setting event-based models often evaluate with the streaming setting [15, 28, 26, 17]. In this setting, information from the previously observed batches of events can be used to update the model however no information from the test set can be used to train the model.
Live-update Setting You et al. [18] proposed the live-update setting where the model weights are constantly updated to newly observed snapshots while predicting the next snapshot. To predict links in $\mathbf{G}_{t+1}$, first the observed snapshot $\mathbf{G}_{t-1}$ are split into a training set and a validation set. The model is trained on $\mathbf{G}_{t-1}^{train}$ while using $\mathbf{G}_{t-1}^{val}$ for early stopping. Lastly, the trained model receives $\mathbf{G}_{t}$ and predicts for $\mathbf{G}_{t+1}$.
@@
Appendix C Temporal Graph Learning Methods
C.1 Snapshot-based Methods
snapshot-based methods receives a sequence of graph snapshots as input, representing the temporal graph at specific time intervals (hours, days, etc.). Therefore, DTDG methods are designed to process entire snapshot at once (often with a graph learning model) and then utilize mechanisms to learn temporal dependencies between snapshots. Example methods are as follows:
  • HTGN. Many DTDG methods focus on learning structural and temporal dependencies in an Euclidean space thus omitting the complex and hierarchical properties which arises in real world networks. To address this, Yang et al. [12] proposed a Hyperbolic Temporal Graph Network (HTGN) which utilizes the exponential capacity and hierarchical awareness of hyperbolic geometric. More specifically, HTGN incorporates hyperbolic graph neural network and hyperbolic gated recurrent neural network to capture the structural and temporal dependencies of a temporal graph, implicitly preserving hierarchical information. In addition, the hyperbolic temporal contextual self-attention module is used to attend to historical states while the hyberbolic temporal consistency module ensures model stability and generalization.
  • GCLSTM. To learn over a sequences of graph snapshots, Chen et al. [13] proposed a novel end-to-end ML model named Graph Convolution Network embedded Long Short-Term Memory (GC-LSTM) for the dynamic link prediction task. In this work, the LSTM act as the main framework to learn temporal dependencies between all snapshots if a temporal graph while GCN is applied on each snapshot to capture the structural dependencies between nodes. Two GCNs are used to learn the hidden state and the cell state for the LSTM and the decoder is a MLP mapping the feature at the current time back to the graph space. The design of GC-LSTM allows it to handle both link additions and link removals.
  • EGCN. Existing approaches often require the knowledge of a node during the entire time span of a temporal graph while real world networks often changes its node set. To address this challenge, Pareja et al. [14] proposed the EvolveGCN (EGCN) model which captures the dynamic of the graph sequence by using an RNN to update the weight of a GCN. In this way, The RNN regulates the GCN model parameter directly and effectively performing model adaptation. This allows node changes because the learning is performed on the model itself, rather than specific sequence of node embeddings. Note that the GCN parameters are not trained and only computed from the RNN. Empirically, the model achieves good performance for link prediction, edge classification and node classification on DTDGs.
  • PyG-Temporal. PyTorch Geometric Temporal (PyG-Temporal) is an open-source Python library which combines state-of-the-art methods for neural spatiotemporal signal processing [22]. Many existing methods such as EGCN, GCLSTM and more are implemented directly in PyG-Temporal for research. PyG-Temporal is designed with a simple and consistent API following existing geometric deep learning library such as Pytorch Geometric [40]. Originally, PyG-Temporal are designed for node level regression tasks on datasets available exclusively within the framework. In this work, we apply PyG-Temporal models for the link prediction tasks on publicly available datasets.
C.2 Event-based Methods
Continuous Time Dynamic Graph (CTDG) methods receive a continuous stream of edges as input and make predictions over any possible timestamps. CTDG methods incorporate newly observed information into its predictions by updating its internal representation of the world. For efficiency, the stream of edges are divided into fixed size batches while predictions are made for each batch sequentially. To incorporate the latest information, edges from each batch becomes available to the model once the predictions are made. Different from DTDG, CTDG has no inherent notion of graph snapshots, models often track internal representations of a node over time and sample temporal neighborhoods surrounding the node of interest for prediction.
  • TGAT. Xu et al. [41] argued that models for temporal graphs should be able to quickly generate embeddings in an inductive fashion when new nodes are encountered. The key component of the proposed Temporal Graph Attention (TGAT) layer is to combine the self-attention mechanism with a novel functional time encoding technique derived from Bochner’s theorem from classical harmonic analysis. In this way, a TGAT layer can efficiently learn from temporal neighborhood features as well as temporal dependencies. The functional time encoding provides a continuous functional mapping from the time domain to a vector space. The hidden vector of time then replaces positional encoding used in the self-attention mechanism.
@@ C.2 Event-based Methods
  • TGN. Rossi et al.. [15] introduce Temporal Graph Network (TGN), a versatile and efficient framework for dynamic graphs, represented as stream of timestamped events. TGN leverage a combination of memory modules and graph-based operators to improve computational efficiency. Essentially, TGN is a framework that subsumes several previous models as specific instances. When making predictions for a new batch, TGN first update the memory with messages coming from previous batches to allow the model to incorporate novel information from observed batches.
@@
  • CAWN. Causal Anonymous Walks (CAWs) [26] are proposed for representing temporal networks inductively to learn the laws governing the link evolution on networks such as the triadic closure law. CAWs, derived from temporal random walks, act as automatic retrievals of temporal network motifs, avoiding the need for their manual selection and counting. An anonymization strategy was also proposed to replace node identities with hitting counts from sampled walks, maintaining inductiveness and motif correlation. CAWN is a neural network model proposed to encode CAWs, paired with a CAW sampling strategy that ensures constant memory and time costs for online training and inference.
@@
  • TCL. TCL [42] effectively learns dynamic node representations by capturing both temporal and topological information. It features three main components: a graph-topology-aware transformer adapted from the vanilla Transformer, a two-stream encoder that independently extracts temporal neighborhood representations of interacting nodes and models their interdependencies using a co-attentional transformer, and an optimization strategy inspired by contrastive learning. This strategy maximizes mutual information between predictive representations of future interaction nodes, enhancing robustness to noise.
@@
  • NAT. In modeling temporal networks, the neighborhood of nodes provides essential structural information for interaction prediction. It is often challenging to extract this information efficiently. Luo et al. [17] propose the Neighborhood-Aware Temporal (NAT) network model that introduces a dictionary-type neighborhood representation for each node. NAT records a down-sampled set of neighboring nodes as keys, enabling fast construction of structural features for joint neighborhoods. A specialized data structure called N-cache is designed to facilitate parallel access and updates on GPUs.
  • EdgeBank. EdgeBank [16] is a non-learnable heuristic baseline which simply memorizes previously observed edges. The surprisingly strong performance of EdgeBank in existing evaluation inspired the authors to also propose novel, more challenging and realistic evaluation protocals for dynamic link prediction.
@@
  • DyGFormer DyGFormer. Yu et al. [28] introduces a transformer-based architecture for dynamic graph learning. DyGFormer focuses on learning from nodes historical first-hop interactions and employs a neighbor co-occurrence encoding scheme to capture correlations between source and destination nodes through their historical sequences. A patching technique was also proposed to divide each sequence into patches for the transformer, enabling effective utilization of longer histories. DyGLib was also presented as a library for standardizing training pipelines, extensible coding interfaces, and thorough evaluation protocols to ensure reproducible dynamic graph learning research.
Appendix D Computing Resources
@@
Appendix E Model Configurations